OpenStudy (anonymous):

how to solve: 3log8+9log2. the 3 and 9 are the base numbers and these must be the same in order to add logarithms

OpenStudy (unklerhaukus):

\[\log_38+\log_92\]

OpenStudy (anonymous):

ow okay, so now both bases are 10

OpenStudy (unklerhaukus):

use the change of base formula on both terms \[\log_b(a) = \frac{\log_{10}(a)}{\log_{10}(b)}\]

OpenStudy (anonymous):

So the answer is log531441 (3^8 x 9^2)

OpenStudy (unklerhaukus):

what ? that dosen't look right, the answer is nice and simple

OpenStudy (unklerhaukus):

\[\log_38+\log_92=\frac{\log8}{\log3}+\frac{\log2}{\log9}\]

OpenStudy (unklerhaukus):

now add the fractions, as you would if they were natural numbers rather than logs: \[\frac ab+\frac cd = \frac{ad+bc}{bd}\] ...

OpenStudy (anonymous):

so \[\frac{ (\log8xlog9)+ \log3 x \log9 }{ \log3 x \log9 }\]

OpenStudy (anonymous):

So the answer must be log72, since you can take away log3 x log9

OpenStudy (unklerhaukus):

yes good work, now lets simplify some of those individual log terms with: \[\log (a^n)=n\log(a)\]

OpenStudy (unklerhaukus):

wahat what did you just say?

OpenStudy (unklerhaukus):

take away?

OpenStudy (unklerhaukus):

wait i think i see a mistake

OpenStudy (anonymous):

the is a log3 x log 9 at the top and the bottom, because of that you can leave them away

OpenStudy (unklerhaukus):

\[\log_38+\log_92=\frac{\log8}{\log3}+\frac{\log2}{\log9}\\ =\frac{ \log8\cdot\log9+ \log3 \cdot \log2 }{ \log3 x \log9 }\]

OpenStudy (anonymous):

ow it must be ac+bc/bd

OpenStudy (unklerhaukus):

\[\frac AB+\frac CD = \frac{AD}{BD}+\frac{BC}{BD}=\frac{AD+BC}{BD}\]

OpenStudy (anonymous):

okay, now I get that. But how can it be simplified?

OpenStudy (unklerhaukus):

can you express 8 like a^n

OpenStudy (anonymous):

Can't you leave the log9 and log 3 away both at the top and bottom so you are left with log8 . log2 ?

OpenStudy (anonymous):

8 is 2^3

OpenStudy (unklerhaukus):

so log 8 = log 2^3 = 3 x log2 what about log 9 = ?

OpenStudy (anonymous):

2. log 3

OpenStudy (unklerhaukus):

We have: \[\log_38+\log_92=\frac{\log8}{\log3}+\frac{\log2}{\log9}\\ =\frac{ \log8\cdot\log9+ \log3 \cdot \log2 }{ \log3 \cdot \log9}\\ =\frac{ (3\times \log2)\cdot(2\times\log3)+ \log3 \cdot \log2 }{ \log3 \cdot( 2\times\log3)}\]

OpenStudy (anonymous):

yes, that's clear

OpenStudy (unklerhaukus):

now simplify the numerator

OpenStudy (anonymous):

\[\frac{ (4 \times \log2)+(3\times \log3 }{\log3 . (2\times \log 3) }\]

OpenStudy (anonymous):

and now you have log2^4+log3^3 at the numerator

OpenStudy (unklerhaukus):

um, not really \[(3\times\log2)\cdot(2\times\log3)=(3\times2)\times \log2\cdot\log3\] so the numerator is \[(3\times2)\times\log2\cdot\log3+\log3\cdot\log2\\ =(3\times2+1)\log_2\log_3\]

OpenStudy (anonymous):

Oh okay I get it now

OpenStudy (unklerhaukus):

typo last line should be^*\[=(3\times2+1)\log(2)\log(3)\]

OpenStudy (anonymous):

So (6 x )log2 x log 3

OpenStudy (anonymous):

sorry 7

OpenStudy (unklerhaukus):

yesh, numerator is 7 log(2) log(3)

OpenStudy (unklerhaukus):

now lets simplify the denominator \[\log3\cdot(2\times\log3)\] use \(\log_b^2(a)\), etc for \(\log_b(a)\times\log_b(a)\)

OpenStudy (anonymous):

log3 . (log3^2) ?

OpenStudy (unklerhaukus):

we want to keep the 2 out the front

OpenStudy (anonymous):

owh so 2 x (log9) ?

OpenStudy (anonymous):

No that's wrong

OpenStudy (unklerhaukus):

log 3 + log 3 is log 9 log 3 x log 3 is not log 9 log 3 x log 3 is log^2 3

OpenStudy (unklerhaukus):

so what does our fraction look like now?

OpenStudy (anonymous):

\[\frac{ 7 \times (2 \times \log3) }{ 2 \times \log 8 }\]

OpenStudy (unklerhaukus):

try again

OpenStudy (anonymous):

\[\frac{ 7 \times (\log2 \times \log3) }{ 2 \times \log3 + \log2^3 }\]

OpenStudy (unklerhaukus):

\[\frac{7\log(2)\log(3)}{2\log^2(3)}=\frac{7\log(2)\log(3)}{2\log(3)\log(3)}\]

OpenStudy (anonymous):

ow because you have log x log= log^2

OpenStudy (unklerhaukus):

don't mix up \[\log(a)+\log(a)=\log(a^2)\]with \[\log(a)\times\log(a)=\log^2(a)\]

OpenStudy (anonymous):

okay didn't know that's possible with log. Clear now

OpenStudy (unklerhaukus):

now we can cancel the common term in the fraction...

OpenStudy (anonymous):

so: \[\frac{ 7 \times \log2 }{ 2 \times \log3 }\]

OpenStudy (anonymous):

canceled log 3 at the top and bottom

OpenStudy (unklerhaukus):

yes,

OpenStudy (unklerhaukus):

now you could leave it like that, with logs to base 10, or use the change of base formula in reverse ... \[\frac{\log(a)}{\log(b)}=\log_b(a)\]

OpenStudy (anonymous):

how you change the base formula in reserve since you're left with the 7 and 2 ?

OpenStudy (unklerhaukus):

take the 7/2 out the front, use the change of base only on log(2) / log(3)

OpenStudy (anonymous):

okay so 7/2 ( 2log3)

OpenStudy (unklerhaukus):

\[\frac72\log_3(2)\] ie (7/2) log_3 (2)

OpenStudy (anonymous):

okay thanks alot :)

OpenStudy (unklerhaukus):

logs are tricky